Work, energy, and energy conservation



1. How much work is done on a 2 kg object when it changes its velocity from 20 m/s to 40 m/s?



2. A pile of books of mass 50 kg is placed on top of a cabinet 3 m high. What is the gravitational


potential energy of the pile of books?



3. A box is pulled by a force of 200 N making an angle of 30 degrees with the horizontal. How


much work is done on the box if it is displaced by 20 m?

1
Expert's answer
2021-11-21T17:27:29-0500

1. "W=\Delta KE=mv^2/2-mv_0^2/2=2\cdot40^2/2-2\cdot20^2/2=1200\ (J)"


2. "PE=mgh=50\cdot9.8\cdot3=1470\ (J)"


3. "W=Fs\cos\alpha=200\cdot20\cdot\cos30°=3464\ (J)"